OpenStudy (arbyj):

how do you turn y=2/3x+4 into standard form?

OpenStudy (tkhunny):

What is the appearance of Standard Form?

OpenStudy (arbyj):

@tkhunny ax+bx=c

OpenStudy (tkhunny):

Have you considered multiplying by 3? Might be a good start.

OpenStudy (shamil98):

in standard form you want whole numbers. which is why you would multiply the equation by 3.

OpenStudy (arbyj):

So nothing becomes negative?

OpenStudy (tkhunny):

It will in a moment. Please show the equation after multiplying by 3.

OpenStudy (arbyj):

2x+y=4, right?

OpenStudy (shamil98):

you multiply everything in the equation by 3 not just 2/3.

OpenStudy (arbyj):

Am I wrong?

OpenStudy (arbyj):

oh yeah

OpenStudy (arbyj):

@shamil98 2x+3y=12, is this right?

OpenStudy (shamil98):

It is -2x + 3y = 12 when you move the x to the other side you subtract it.

OpenStudy (arbyj):

Okay thank you!

OpenStudy (tkhunny):

Some versions want a > o.